OpenStudy (anonymous):

After a knee surgery, your trainer tells you to return to your jogging program slowly. He suggests jogging for 12 minutes each day for the first week. Each day thereafter, he suggests that you increase that time by 6 minutes per day. How many weeks will it be before you are up to jogging 60 minutes per day?

OpenStudy (anonymous):

You complain that the hot tub in your hotel suite is not hot enough. The hotel tells you that they will increase the temperature by 10% each hour. If the current temperature of the hot tub is 75 degrees F, what will be the temperature of the hot tub after 3 hours, to the nearest tenth of a degree?

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

Do i use sum formula, seriies?

OpenStudy (aripotta):

i haven't learned it yet :/

OpenStudy (anonymous):

its cool i think its 9 weeks lol, just figured it out, but do you know the second part of the question by any chance?

OpenStudy (aripotta):

99.825?

OpenStudy (aripotta):

oh, nearest tenth degree would be 99.8

OpenStudy (anonymous):

How did you get that, please explain

OpenStudy (aripotta):

75*0.1 = 7.5. 7.5 + 75 = 82.5 (10% after first hour) 82.5*0.1 = 8,25. 8.25 + 82.5 = 90.75 (10% after second hour) 90.75*0.1 = 9.075. 9.075 + 90.75 = 99.825 (10% after third hour)

OpenStudy (anonymous):

i get how you did 7.5 *.01 bu then you did 7.5.7.5, multiplied?

OpenStudy (anonymous):

oh never mind, thats a period

OpenStudy (aripotta):

yea, sorry :/

OpenStudy (anonymous):

Its alright, thanks for the help though, your great :D

OpenStudy (aripotta):

:3